# InterCLIP-MEP: Interactive CLIP and Memory-Enhanced Predictor for Multi-modal Sarcasm Detection

### v5

@@
Abstract
Sarcasm in social media, often expressed through text-image combinations, poses challenges for sentiment analysis and intention mining. Current multi-modal sarcasm detection methods have been demonstrated to overly rely on spurious cues within the textual modality, revealing a limited ability to genuinely identify sarcasm through nuanced text-image interactions. To solve this problem, we propose InterCLIP-MEP, which introduces Interactive CLIP (InterCLIP) with an efficient training strategy to extract enriched text-image representations by embedding cross-modal information directly into each encoder. Additionally, we design a Memory-Enhanced Predictor (MEP) with a dynamic dual-channel memory that stores valuable test sample knowledge during inference, acting as a non-parametric classifier for robust sarcasm recognition. Experiments on two benchmarks demonstrate that InterCLIP-MEP achieves state-of-the-art performance, with significant accuracy and F1 score improvements on MMSD and MMSD2.0. Our code is available at https://github.com/CoderChen01/InterCLIP-MEP.
@@ Abstract
InterCLIP-MEP: Interactive CLIP and Memory-Enhanced Predictor for Multi-modal Sarcasm Detection
Junjie Chen1,3thanks: Junjie is a research assistant at the EPIC Lab of Shanghai Jiao Tong University, working remotely with Linfeng Zhang.    Hang Yu2thanks: Co-corresponding author.    Subin Huang3thanks: Corresponding author.    Sanmin Liu3    Linfeng Zhang1 1Shanghai Jiao Tong University, China    2Shanghai University, China 3Anhui Polytechnic University, China jorji.chen@gmail.com,  yuhang@shu.edu.cn,  subinhuang@ahpu.edu.cn, sanmin.liu@ahpu.edu.cn,  zhanglinfeng@sjtu.edu.cn
1 Introduction
Sarcasm, with its subtlety and complexity, plays a key role in communication by conveying irony, mockery, or hidden meanings Muecke (1982); Gibbs and O’Brien (1991); Gibbs and Colston (2007). Automatically detecting sarcasm in text has become an important research area, supporting tasks like sentiment analysis and intent mining Pang et al. (2008); Tsur et al. (2010); Bouazizi and Ohtsuki (2015).
Figure 1: An overview of the shortcomings of existing multi-modal sarcasm detection pipelines. In panels (a) and (b), we present two main multi-modal sarcasm detection pipelines, with shortcomings indicated by a red question mark. In panel (c), we visually show an example of multi-modal sarcasm cues correctly or incorrectly recognized in a multi-modal sarcasm sample.
With the rise of social media platforms like Twitter and Reddit, users often use text-image combinations to express their messages. As a result, multi-modal sarcasm detection is increasingly important, posing challenges in understanding the complex relationship between textual and visual cues to identify sarcasm.
@@ 1 Introduction
As shown in Figures 1(a) and 1(b), many methods rely on dual unimodal pre-trained encoders, such as ViTDosovitskiy et al. (2021) and BERT Devlin et al. (2019), as the backbone for encoding text-image pairs, followed by specific feature fusion Xu et al. (2020); Pan et al. (2020); Liang et al. (2021, 2022); Wen et al. (2023); Tian et al. (2023); Wei et al. (2024). However, this approach may not capture multi-modal sarcasm cues as effectively as multi-modal pre-trained models like CLIP Radford et al. (2021). In Figure 1(a), the use of a learnable classification head to predict labels from fused representations is common but often associated with high predictive entropy and significant uncertainty. Wei et al. (2024) pioneered the construction of a static semantic space using historical training samples, where more robust predictions are obtained during inference through KNN-based querying and voting, as illustrated in Figure 1(b). However, while CLIP has been proven to be effective in serving as a text-image encoder for multi-modal sarcasm detection Qin et al. (2023), it still struggles to capture multi-modal sarcasm cues due to the inherent inconsistency of sarcasm, which conflicts with CLIP’s direct alignment of text and image. Furthermore, relying on a static semantic space for inference is ill-suited to handle the dynamic nature of evolving sample distributions. In fact, Qin et al. (2023) have shown that many models rely on spurious cues in the MMSD benchmark Cai et al. (2019), leading to biased results.
Building on the limitations of prior multi-modal sarcasm detection approaches, we propose Interactive CLIP (InterCLIP) as the backbone, embedding cross-modal representations directly into text and vision encoders to enhance the understanding of multi-modal sarcasm cues (Figure 2, left). To complement this, we design a Memory-Enhanced Predictor (MEP) that dynamically utilizes historical test sample features to create a more adaptive and reliable non-parametric classifier for final predictions (Figure 2, right). Together, these components form the proposed framework, InterCLIP-MEP. Furthermore, InterCLIP-MEP employs an efficient training strategy that fine-tunes cross-modal interactions through a lightweight adaptation mechanism, ensuring computational efficiency while delivering state-of-the-art performance (Figure 2, left). Overall, our contributions are as follows:
@@
  • We introduce InterCLIP-MEP [^1] Our code is available in the supplementary material., a novel framework for multi-modal sarcasm detection, which combines Interactive CLIP (InterCLIP) for enhanced text-image interaction encoding and Memory-Enhanced Predictor (MEP) for more robust and reliable sarcasm predictions.
  • We propose an efficient training strategy that significantly reduces computational overhead compared to state-of-the-art methods. By introducing approximately 20.6x fewer trainable parameters, our approach reduces GPU memory usage by about 2.5x and accelerates computation by roughly 8.7x with a batch size of 128, all while maintaining superior performance on a single NVIDIA RTX 4090 GPU.
  • Through extensive experiments on the MMSD and MMSD2.0 benchmarks, we show that InterCLIP-MEP improves accuracy by 1.08% and F1 score by 1.51% over state-of-the-art methods, especially on MMSD2.0.
2 Related Work
Early research in sarcasm detection focused primarily on text data Bouazizi and Ohtsuki (2015); Amir et al. (2016); Baziotis et al. (2018). With the rise of social media, detecting sarcasm in text-image pairs has become more challenging, driving the development of multi-modal approaches. Schifanella et al. (2016) were among the first to explore multi-modal social media posts for identifying sarcasm cues. Building on this, Cai et al. (2019) introduced the MMSD benchmark, demonstrating the effectiveness of a hierarchical fusion model that integrates image features. This benchmark has since become a foundation for multi-modal sarcasm detection, inspiring subsequent studies Xu et al. (2020); Pan et al. (2020); Liang et al. (2021, 2022); Liu et al. (2022); Qin et al. (2023); Wen et al. (2023); Tian et al. (2023); Wei et al. (2024).
However, the MMSD benchmark was later found to contain spurious cues that could lead to model bias Qin et al. (2023). To mitigate this, Qin et al. (2023) introduced the MMSD2.0 benchmark, which removes these cues and corrects mislabeled samples. Re-evaluations on MMSD2.0 revealed significant performance drops in existing models, emphasizing the need for more robust approaches. In parallel, Tang et al. (2024) explored the use of large language models (LLMs) in multi-modal sarcasm detection, incorporating instruction templates and retrieval modules. While promising, the performance improvements were modest compared to the substantial computational cost.
In this work, we present InterCLIP-MEP, a lightweight and efficient framework that achieves competitive performance without the high resource demands of LLM-based approaches. By overcoming the limitations of current methods, our approach offers a practical and scalable solution for multi-modal sarcasm detection.
@@ 2 Related Work
Figure 2: Overview of our framework. (I) Training Interactive CLIP (InterCLIP): Vision and text representations are extracted using separate encoders and embedded into the top-$n$ layers of the opposite modality’s encoder for interaction. The top-$n$ layers are fine-tuned with LoRA, while the rest of the encoder remains frozen. Final vision and text representations are concatenated and used to train a classification module for identifying multi-modal sarcasm. A projection module is also trained to project representations into a latent space. (II) Memory-Enhanced Predictor (MEP): During inference, InterCLIP generates interactive representations. The classification module assigns pseudo-labels, and the projection module provides projection features. MEP updates dynamic memory with these features and pseudo-labels. The final prediction of the current sample is made by comparing its projected feature with those in memory.
3 Methodology
An overview of InterCLIP-MEP is illustrated in Figure 2. Initially, we elaborate on the Interactive CLIP (InterCLIP) and its training strategy, followed by an in-depth explanation of the Memory-Enhanced Predictor (MEP).
3.1 Interactive CLIP
@@
The input to Interactive CLIP (InterCLIP) is a text-image pair $\mathcal{P}=(T,I)$, where $T$ represents a piece of text and $I$ represents an image. Here, for simplicity, we do not consider the case of batch inputs.
The text encoder $\mathcal{T}$ extracts the vanilla text representations $\mathbf{F}_{t}$:
$\mathbf{F}_{t}$ $=\mathcal{T}(T)$
=$=\{h^{t}_{\textrm{bos}}(t_{\textrm{bos}}),h^{t}_{1}(t_{1}),\dots,h^{t}_{n}(t_{n}),h^{t}_{\textrm{eos}}(t_{\textrm{eos}})\},$ (1)
@@
$\mathbf{F}_{v}=\mathcal{V}(I)=\{h^{v}_{\textrm{cls}}(p_{\textrm{cls}}),h^{v}_{1}(p_{1}),\dots,h^{v}_{m}(p_{m})\},$ (2)
@@
$\tilde{\mathbf{F}}_{t}=\mathcal{T}(T|\mathbf{F}_{v}),\tilde{\mathbf{F}}_{v}=\mathcal{V}(V|\mathbf{F}_{t}).$ (3)
@@
To fuse the input representations $\mathbf{H}_{t/v}$ with the projected embedded representations $\mathbf{F}_{v/t}^{{}^{\prime}}=\mathcal{F}_{t/v}(\mathbf{F}_{v/t})$, we concatenate them and feed them into the attention layer to obtain the transformed representations. We then extract the transformed input representations $\mathbf{H}^{{}^{\prime}}_{t/v}$ from the output. Following Ganz et al. (2024), we apply a gated projection layer $\mathcal{G}_{t/v}$ along with the self-attention’s projection head $\mathcal{H}_{t/v}$ using a learnable gating mechanism to compute the self-attention output representations $\mathbf{H}^{{}^{\prime\prime}}_{t/v}$. Given the similarity between the self-attention layers of the vision encoder and the text encoder, we use the text encoder $\mathcal{T}$ to illustrate the process as follows:
$\begin{split}\mathbf{F}_{v}^{{}^{\prime}}&=\mathcal{F}_{t}(\mathbf{F}_{v}),\quad\mathbf{F}_{v}\in\mathbb{R}^{m\times d_{v}},\mathbf{F}_{v}^{{}^{\prime}}\in\mathbb{R}^{m\times d_{t}},\\ \mathbf{H}^{{}^{\prime}}_{t}&=\textrm{Attn}_{t}(\mathbf{H}_{t}\oplus\mathbf{F}_{v}^{{}^{\prime}})_{[:n]},\\ &\phantom{=}\mathbf{H}_{t},\mathbf{H}^{{}^{\prime}}_{t}\in\mathbb{R}^{n\times d_{t}},\mathbf{H}_{t}\oplus\mathbf{F}_{v}^{{}^{\prime}}\in\mathbb{R}^{(n+m)\times d_{t}},\\ \mathbf{H}^{{}^{\prime\prime}}_{t}&=\mathcal{H}_{t}(\mathbf{H}^{{}^{\prime}}_{t})+\mathcal{G}_{t}(\mathbf{H}^{{}^{\prime}}_{t})\cdot\textrm{tanh}(\beta_{t}),\mathbf{H}^{{}^{\prime\prime}}_{t}\in\mathbb{R}^{n\times d_{t}}.\end{split}$ (4)
Here, $\oplus$ denotes the concatenation operation, and $\beta_{t}$ is a learnable gating parameter initialized to 0 to ensure training stability. The subsequent computation follows the original CLIP Radford et al. (2021), ultimately yielding the interactive representations $\tilde{\mathbf{F}}_{t}$.
Figure 3: Structure of the conditional self-attention.

### v7

@@
Abstract.
[…]
@@ Abstract.
Multi-modal Sarcasm Detection, Multi-modal Interaction, Cross-modal Representations, Memory-enhanced Predictor, Parameter-efficient Fine-tuning
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXjournal: JACMjournalvolume: 37journalnumber: 4article: 111publicationmonth: 8ccs: Information systems Multimedia information systemsccs: Computing methodologies Information extractionccs: Information systems Sentiment analysis
1. Introduction
Sarcasm, characterized by its nuanced and intricate nature, serves as a vital mechanism in communication, often employed to convey irony, mockery, or implicit meanings (Gibbs1991PsychologicalAO; tsur2010icwsm; JingSOJN23; LiuZS24). The automated identification of sarcasm in textual data has evolved into a significant research focus, offering substantial support for tasks such as sentiment analysis and intent mining (pang2008opinion; JoshiTPBC16; JoshiBC17). With the growing prevalence of social media platforms such as Twitter and Reddit, users increasingly rely on the fusion of textual and visual elements to articulate their messages. Consequently, multi-modal sarcasm detection has gained considerable prominence, presenting challenges in deciphering the intricate interplay between textual and visual cues to accurately identify sarcastic content (WangYJMXL24; ZhuangLZGZM25).
Despite significant progress achieved by prior studies (xu2020reasoning; pan-etal-2020-modeling; liang2021multi; Liang2022MultiModalSD; 10.24963/ijcai.2024/887) in this domain, Qin et al. (qin2023mmsd2) have identified the presence of spurious cues within the MMSD benchmark (Cai2019MultiModalSD) utilized in these works. These cues contribute to the development of biased models, thereby overestimating their effectiveness in detecting sarcasm cues from multi-modal data. Through the introduction of MMSD2.0, a refined benchmark that eliminates spurious cues and rectifies mislabeled samples, Qin et al. (qin2023mmsd2) reveal a significant performance degradation of existing methods when evaluated on MMSD2.0. Although their proposed method, which leverages Multi-view CLIP for reliable multi-modal sarcasm detection, shows promise, it still exhibits limited capability in accurately discerning sarcasm through subtle text-image interactions.
Figure 1. An overview of the limitations inherent in existing multi-modal sarcasm detection frameworks. Panels (a) and (b) illustrate two predominant multi-modal sarcasm detection pipelines, with their respective shortcomings highlighted by red question marks. Panel (c) provides a visual representation of multi-modal sarcasm cues, demonstrating instances where such cues are either accurately identified or misinterpreted in a multi-modal sarcasm sample.
@@ 1. Introduction
As depicted in Figures 1(a) and 1(b), numerous approaches utilize dual unimodal pre-trained encoders, such as ViT (dosovitskiy2020image) and BERT (devlin2019bert), as the backbone for encoding text-image pairs, followed by specialized feature fusion techniques (xu2020reasoning; pan-etal-2020-modeling; liang2021multi; Liang2022MultiModalSD; wen2023dip; Tian2023DynamicRT; wei2024g2sam; Wang2025AdaSFFuse). However, such methodologies may fail to effectively capture multi-modal sarcasm cues compared to multi-modal pre-trained models like CLIP (radford2021learning). In Figure 1(a), the adoption of a learnable classification head to predict labels from fused representations is prevalent but often accompanied by high predictive entropy and substantial uncertainty. The work by (wei2024g2sam) introduced the construction of a static semantic space utilizing historical training samples, where robust predictions during inference are achieved through KNN-based querying and voting, as illustrated in Figure 1(b). Nevertheless, despite CLIP’s demonstrated effectiveness as a text-image encoder for multi-modal sarcasm detection (qin2023mmsd2), it remains challenged in capturing multi-modal sarcasm cues due to the inconsistency of sarcasm, which contradicts CLIP’s alignment of text and image. Additionally, reliance on a static semantic space for inference is inadequate to address the dynamic nature of evolving sample distributions. Indeed, Qin et al. (qin2023mmsd2) have highlighted that many models depend on spurious cues within the MMSD benchmark (Cai2019MultiModalSD), resulting in biased outcomes.
In this paper, building on these insights, we propose InterCLIP-MEP, a novel framework designed to enhance the reliability of multi-modal sarcasm detection. Our InterCLIP-MEP introduces a refined variant of CLIP, Interactive CLIP (InterCLIP), as the backbone to enhance sample encoding for detecting multi-modal sarcasm. InterCLIP directly embeds cross-modal representations into text and vision encoders, thereby enhancing the understanding of multi-modal sarcasm cues (as illustrated in Figure 2, on the left). To complement this, we further design a Memory-Enhanced Predictor (MEP), which dynamically leverages historical test sample features to construct a more adaptive and reliable non-parametric classifier (Figure 2, right). Importantly, the design of MEP is conceptually motivated by both cognitive science and machine learning theories. From the cognitive perspective, human working memory is known to have limited capacity and tends to preferentially retain information with higher certainty (Baddeley2000TheEB; anderson1991reflections). Inspired by this mechanism, MEP employs an entropy-based update strategy to selectively preserve low-uncertainty features, simulating human-like selective retention. From the machine learning perspective, entropy is a widely recognized measure of predictive uncertainty, and uncertainty-aware sample selection has been shown effective in active learning and memory-based inference (toneva2018empirical; settles2009active). Beyond these general motivations, the characteristics of sarcasm further justify our design: sarcastic expressions in social media typically fall into a limited set of semantic patterns (e.g., exaggeration, contrast, or ironic praise), and humans often rely on prior experiences or historical contexts to correctly interpret them (attardo2000irony; gibbs1986psycholinguistics; gibbs2007irony; bamman2015contextualized; wallace-etal-2014-humans). By maintaining low-entropy representations in memory, MEP effectively acts as a denoising mechanism, ensuring more robust predictions during inference. Collectively, these components constitute the proposed framework, InterCLIP-MEP. Additionally, InterCLIP-MEP incorporates an efficient training strategy that fine-tunes cross-modal interactions through a lightweight adaptation mechanism, ensuring computational efficiency while achieving state-of-the-art performance (Figure 2, left).
In summary, the key contributions of this work are as follows:
@@
  • The proposed framework, InterCLIP-MEP, introduces a novel approach for multi-modal sarcasm detection. It integrates Interactive CLIP (InterCLIP) to enhance text-image interaction encoding and incorporates a Memory-Enhanced Predictor (MEP) to achieve more robust and reliable sarcasm predictions.
[…]
  • Extensive experiments conducted on the MMSD, MMSD2.0 and DocMSU benchmarks demonstrate that InterCLIP-MEP achieves significant improvements, outperforming state-of-the-art methods with an increase of 1.08% in accuracy and 1.51% in F1 score, particularly on the MMSD2.0 dataset.
The remainder of this paper is structured as follows. Section 2 reviews related studies and discusses the limitations of existing methodologies. Section 3 introduces the proposed InterCLIP-MEP framework, including the Interactive CLIP backbone, efficient training strategy, and Memory-Enhanced Predictor. Section 4 presents the main experiments, which provide both quantitative and qualitative evaluations as well as an efficiency analysis to demonstrate the effectiveness of the proposed method. Section 5 reports the extended experiments, covering four aspects: performance on a document-level benchmark, evaluation under low-resource settings, error propagation robustness study, and robustness analysis under distributional shifts, to comprehensively verify the generalization and stability of InterCLIP-MEP. Section 6 discusses the limitations of the current study and outlines potential directions for future research. Finally, Section 7 concludes the paper.
2. Related Work
2.1. Multi-modal Sarcasm Detection
Early research in sarcasm detection predominantly focused on textual data (bouazizi2015sarcasm; amir2016modelling; baziotis2018ntua). With the proliferation of social media platforms, identifying sarcasm in text-image pairs has emerged as a more complex challenge, prompting the advancement of multi-modal methodologies. Among the pioneering efforts, Schifanella et al. (schifanella2016detecting) explored multi-modal social media posts to detect sarcasm cues. Expanding on this work, Cai et al. (Cai2019MultiModalSD) proposed the MMSD benchmark, showcasing the efficacy of a hierarchical fusion model that incorporates image features. This benchmark has since served as a cornerstone for multi-modal sarcasm detection, motivating a series of subsequent studies (xu2020reasoning; pan-etal-2020-modeling; liang2021multi; Liang2022MultiModalSD; Liu2022TowardsMS; qin2023mmsd2; wen2023dip; Tian2023DynamicRT; wei2024g2sam; 10.24963/ijcai.2024/887; ding2022multi; 10477507; zhuang2025multi). However, subsequent studies revealed that the MMSD benchmark contained spurious cues, which could introduce bias into model predictions (qin2023mmsd2). To address this issue, Qin et al. (qin2023mmsd2) developed the MMSD2.0 benchmark, which eliminates these cues and rectifies mislabeled samples. Re-evaluations conducted on MMSD2.0 demonstrated significant performance declines in existing models, underscoring the necessity for more robust methodologies.
[…]
2.2. CLIP Adaptation
The Contrastive Language-Image Pretraining (CLIP) model (radford2021learning) exhibits remarkable capabilities in cross-modal alignment for vision-language understanding tasks. Recent progress in domain-specific adaptations of CLIP has led to substantial performance enhancements across diverse applications, including phrase localization (li2022adapting), open-vocabulary semantic segmentation (Liang2023Open), and action recognition (wang2023seeing). These achievements highlight CLIP’s adaptability in addressing specialized visual-linguistic challenges. Closely related to our work, Ganz et al. (ganz2024question) proposed the Question-Aware Vision Transformer (QA-ViT), which augments frozen vision encoders with question embeddings to produce question-aware visual representations. However, QA-ViT was designed for general multi-modal reasoning tasks such as visual question answering rather than sarcasm detection. It only explores the text-to-vision direction, leaving the reverse integration of visual cues into textual encoding unexplored. More importantly, it does not address the pragmatic and incongruity-based nature of sarcasm, where subtle mismatches between modalities must be captured.
In parallel, Qin et al. (qin2023mmsd2) introduced Multi-view CLIP as the first CLIP-based approach for multi-modal sarcasm detection, where frozen CLIP encoders are used to extract text-only, image-only, and joint representations, and the outputs are fused via a transformer. While pioneering, this design treats CLIP merely as a static feature extractor and lacks adaptive cross-modal interaction within the encoders. In contrast, our Interactive CLIP (InterCLIP) explicitly conditions both encoders on cross-modal signals through bidirectional interaction and lightweight LoRA tuning, enabling a richer alignment of subtle sarcastic cues. When combined with our Memory-Enhanced Predictor (MEP), which adaptively incorporates reliable exemplars at inference time, our framework directly targets the unique challenges of sarcasm detection and provides a principled, task-specific advancement over both QA-ViT and Multi-view CLIP.
2.3. Memory-enhanced Prediction
Inspired by the insights from cognitive science (stokes2015activity; Baddeley2000TheEB; anderson1991reflections; gibbs2007irony), the concept of memory has been ingeniously integrated into neural networks to enhance their capabilities (weston2014memory; sukhbaatar2015end). This integration has opened up new avenues for improving model training, as evidenced by several pioneering studies (wu2018unsupervised; wen2023dip; zhu2020label). For instance, Wen et al. (wen2023dip) proposed DIP, which maintains semantic distributions in memory banks to stabilize training and better model incongruity, while Zhu and Yang (zhu2020label) introduced a label-independent memory (LIM) to cache unlabeled video features for robust prototype construction in semi-supervised few-shot learning. Despite differences in domains, both DIP and LIM adopt training-time memory to act as stabilizers or prototype aggregators, and their memories vanish once the model enters the inference stage. In contrast, our Memory-Enhanced Predictor (MEP) introduces a fundamentally different usage of memory: it is a test-time, entropy-aware predictor rather than a training aid. Instead of serving as a regularizer during optimization, MEP dynamically maintains a compact buffer of reliable test exemplars and integrates them into the prediction process through cosine matching. This makes memory a core inference component rather than an auxiliary training tool.
Furthermore, some recent works (zhang2024dual; wei2024g2sam) have also explored inference-time memories. Zhang et al. (zhang2024dual) proposed Dual Memory Networks for general vision-language adaptation, where a static memory caches training features and a dynamic memory accumulates test representations, with cross-attention used to derive adaptive classifiers. Wei et al. (wei2024g2sam) introduced G2SAM for multi-modal sarcasm detection, which leverages a graph-based memory to capture global semantic relations. However, both approaches rely on static or pre-constructed memory structures that do not evolve with incoming test samples. By contrast, our MEP employs a dynamic, entropy-aware buffer that continuously updates with reliable test exemplars, making memory an adaptive and lightweight inference component tailored specifically for the challenges of sarcasm detection (gibbs2007irony; bamman2015contextualized).
Table 1. List of symbols.
Symbol Description
$T$ $T$ denotes a short text.
II II represents an image.
$\mathcal{P}$ $\mathcal{P}$ denotes a text-image pair $(T,I)$.
$\mathcal{T}$ $\mathcal{T}$ denotes CLIP’s text encoder.
$\mathcal{V}$ $\mathcal{V}$ denotes CLIP’s vision encoder.
3. Methodology
An overview of InterCLIP-MEP is illustrated in Figure 2. The framework is first introduced through a detailed exposition of Interactive CLIP (InterCLIP) and its associated efficient training strategy, followed by a comprehensive analysis of the Memory-Enhanced Predictor (MEP). For clarity, Table 1 provides a summary of the primary mathematical symbols employed throughout this work.
@@ 3. Methodology
Figure 2. Overview of our framework. (I) Training Interactive CLIP (InterCLIP): Vision and text representations are extracted using separate encoders and embedded into the top-$n$ layers of the opposite modality’s encoder for interaction. The top-$n$ layers are fine-tuned with LoRA, while the rest of the encoder remains frozen. Final vision and text representations are concatenated and used to train a classification module for identifying multi-modal sarcasm. A projection module is also trained to project representations into a latent space. (II) Memory-Enhanced Predictor (MEP): During inference, InterCLIP generates interactive representations. The classification module assigns pseudo-labels, and the projection module provides projection features. MEP updates dynamic dual-channel memory with these features and pseudo-labels. The final prediction of the current sample is made by comparing its projected feature with those in memory.
3.1. Interactive CLIP
@@
The input to Interactive CLIP (InterCLIP) is a text-image pair $\mathcal{P}=(T,I)$, where $T$ represents a piece of text and $I$ represents an image. Here, for simplicity, we do not consider the case of batch inputs. The text encoder $\mathcal{T}$ extracts the vanilla text representations $\mathbf{F}_{t}$:
(1) $\mathbf{F}_{t}=\mathcal{T}(T)=\{h^{t}_{\textrm{bos}}(t_{\textrm{bos}}),h^{t}_{1}(t_{1}),\dots,h^{t}_{n}(t_{n}),h^{t}_{\textrm{eos}}(t_{\textrm{eos}})\},$
@@
(2) $\mathbf{F}_{v}=\mathcal{V}(I)=\{h^{v}_{\textrm{cls}}(p_{\textrm{cls}}),h^{v}_{1}(p_{1}),\dots,h^{v}_{m}(p_{m})\},$
@@
(3) $\tilde{\mathbf{F}}_{t}=\mathcal{T}(T|\mathbf{F}_{v}),\tilde{\mathbf{F}}_{v}=\mathcal{V}(V|\mathbf{F}_{t}).$
Figure 3. Structure of the conditional self-attention.
@@
To fuse the input representations $\mathbf{H}_{t/v}$ with the projected embedded representations $\mathbf{F}_{v/t}^{{}^{\prime}}=\mathcal{F}_{t/v}(\mathbf{F}_{v/t})$, we concatenate them and feed them into the attention layer to obtain the transformed representations. We then extract the transformed input representations $\mathbf{H}^{{}^{\prime}}_{t/v}$ from the output. Following (ganz2024question), we apply a gated projection layer $\mathcal{G}_{t/v}$ along with the self-attention’s projection head $\mathcal{H}_{t/v}$ using a learnable gating mechanism to compute the self-attention output representations $\mathbf{H}^{{}^{\prime\prime}}_{t/v}$. Given the similarity between the self-attention layers of the vision encoder and the text encoder, we use the text encoder $\mathcal{T}$ to illustrate the process as follows:
(4) $\mathbf{F}_{v}^{{}^{\prime}}=\mathcal{F}_{t}(\mathbf{F}_{v}),\quad\mathbf{F}_{v}\in\mathbb{R}^{m\times d_{v}},\mathbf{F}_{v}^{{}^{\prime}}\in\mathbb{R}^{m\times d_{t}},$
(5) $\mathbf{H}^{{}^{\prime}}_{t}=\textrm{Attn}_{t}(\mathbf{H}_{t}\oplus\mathbf{F}_{v}^{{}^{\prime}})_{[:n]},\quad\mathbf{H}_{t},\mathbf{H}^{{}^{\prime}}_{t}\in\mathbb{R}^{n\times d_{t}},\mathbf{H}_{t}\oplus\mathbf{F}_{v}^{{}^{\prime}}\in\mathbb{R}^{(n+m)\times d_{t}},$
(6) $\mathbf{H}^{{}^{\prime\prime}}_{t}=\mathcal{H}_{t}(\mathbf{H}^{{}^{\prime}}_{t})+\mathcal{G}_{t}(\mathbf{H}^{{}^{\prime}}_{t})\cdot\textrm{tanh}(\beta_{t}),\mathbf{H}^{{}^{\prime\prime}}_{t}\in\mathbb{R}^{n\times d_{t}}.$
Here, $\oplus$ denotes the concatenation operation, and $\beta_{t}$ is a learnable gating parameter initialized to 0 to ensure training stability. The subsequent computation follows the original CLIP, ultimately yielding the interactive representations $\tilde{\mathbf{F}}_{t}$.